OpenStudy (anonymous):

what are the points where a vertical and a horizontal tangenet line occur on the curve x+y^2-y=1. If none, explain. I used implicit differentiation to get 1/1-2y. where do i go from there?

OpenStudy (amistre64):

a vertical line is defined when dx/dy = 0 a horizontal line is defined when dy/dx = 0

OpenStudy (amistre64):

x+y^2-y=1 x' + 2yy' -y' = 0 dx/dy: x' + 2y -1 = 0 dy/dx: 1 + y'(2y-1) = 0

OpenStudy (amistre64):

2y-1 = 0 and 1/(1-2y) = 0 this might not work out

OpenStudy (amistre64):

hmm x = 1+y -y^2 x' = 1 - 2y = 0 when y=1/2 pfft, we have a parabola on its side, no wonder theres no horizontal line